# A polynomial bound on the number of minimal separators and potential maximal cliques in $P_{6}$-free graphs of bounded clique number

### v2

@@ 1 Introduction
Bouchitté and Todinca [2] (with some results generalized by Fomin, Todinca, and Villanger [7]) showed that indeed these notions are useful to solve a wide family of graph problems.
Theorem 1 ([2, 3]).
@@
Theorem 2 (informal statement of [2, 7]).
@@
As a result, if for a graph class $\mathcal{G}$ one can prove a (purely graph-theoretical) polynomial bound on the number of minimal separators or potential maximal cliques in $\mathcal{G}$, then one can immediately obtain polynomial-time algorithms for a wide family of problems on $\mathcal{G}$. We call such a graph class tame. Recently, a methodological study of which graph classes are tame was initiated [9, 10, 1, 14, 8].
@@
We remark that some additional condition to just $P_{6}$-freeness is needed for the conclusion of Theorem 3, as even the class of $P_{5}$-free graphs is not tame. This can be witnessed by prisms: An $n$-prism consists of two $n$-vertex cliques with a matching in between; it is $P_{5}$-free but admits $22^{n}-$ 2 minimal separators. Furthermore, the $P_{6}$ cannot be replaced with $P_{7}$: [5] provides a construction of $P_{7}$-free graphs $G_{n}$ that have clique number $2$, $|V(G_{n})|=6n+2$, and at least $3^{n}$ minimal separators.
Our motivation for proving Theorem 3 is two-fold. First, the proof is very simple, much simpler than the arguments giving polynomial-time algorithms for Maximum Weight Independent Set and related problems in $P_{6}$-free graphs without any assumption on the clique number [11]. Second, it gives examples of graph classes that are tame, but in which such problems as $5$-Coloring or Odd Cycle Transversal are NP-hard: $5$-Coloring is NP-hard in $P_{6}$-free graphs [13] (and trivial in graphs of clique number larger than $5$) and Odd Cycle Transversal is NP-hard in $P_{6}$-free graphs of clique number at most $3$ [6, 4]. This answers negatively a question of Milanič and Pivač [14, Open problem 3].
@@ 2 Proof of Theorem 3
We need the following properties of modules, which can be distilled from Lemma 2 and Theorem 2 and the discussion around them in [12].
@@
We will also need the following lemma of [11].
Lemma 7 (cf. Lemma 4.2 of [11]).
@@
Acknowledgement.
We are grateful to Jakub Gajarský, Lars Jaffke, Paloma Thomé de Lima, Jana Masařiková, Uéverton Souza, and Michał Pilipczuk for inspiring discussions and to Martin Milanič for bringing our attention to [14, Open problem 3].

### v3

@@ Abstract
keywords:
potential maximal cliques, minimal separators, tame class, $P_{6}$-free graphs
@@ 1 Introduction
Bouchitté and Todinca (2001) (with some results generalized by Fomin et al. (2015)) showed that indeed these notions are useful to solve a wide family of graph problems.
Theorem 1 (Bouchitté and Todinca (2001, 2002)).
@@
Theorem 2 (Bouchitté and Todinca (2001); Fomin et al. (2015), informal statement).
@@
As a result, if for a graph class $\mathcal{G}$ one can prove a (purely graph-theoretical) polynomial bound on the number of minimal separators or potential maximal cliques in $\mathcal{G}$, then one can immediately obtain polynomial-time algorithms for a wide family of problems on $\mathcal{G}$. We call such a graph class tame. Recently, a methodological study of which graph classes are tame was initiated, see, e.g., the work of Gartland and Lokshtanov (2023a, b); Abrishami et al. (2022); Milanič and Pivač (2021); Gajarský et al. (2022).
@@
We remark that some additional condition to just $P_{6}$-freeness is needed for the conclusion of Theorem 3, as even the class of $P_{5}$-free graphs is not tame. This can be witnessed by prisms: An $n$-prism consists of two $n$-vertex cliques with a matching in between; it is $P_{5}$-free but admits $22^{n}-$ 2 minimal separators. Furthermore, the $P_{6}$ cannot be replaced with $P_{7}$: Chudnovsky et al. (2023) provide a construction of $P_{7}$-free graphs $G_{n}$ that have clique number $2$, $|V(G_{n})|=6n+2$, and at least $3^{n}$ minimal separators.
Our motivation for proving Theorem 3 is two-fold. First, the proof is very simple, much simpler than the arguments of Grzesik et al. (2022) giving polynomial-time algorithms for Maximum Weight Independent Set and related problems in $P_{6}$-free graphs without any assumption on the clique number. Second, it gives examples of graph classes that are tame, but in which such problems as $5$-Coloring or Odd Cycle Transversal are NP-hard: Huang (2016) proved that $5$-Coloring is NP-hard in $P_{6}$-free graphs (and trivial in graphs of clique number larger than $5$) and Dabrowski et al. (2020); Chudnovsky et al. (2021) proved that Odd Cycle Transversal is NP-hard in $P_{6}$-free graphs of clique number at most $3$. This answers negatively a question of (Milanič and Pivač, 2021, Open problem 3).
@@ 2 Proof of Theorem 3
We need the following properties of modules, which can be distilled from Lemma 2 and Theorem 2 and the discussion around them in the work of Habib and Paul (2010).
@@
We will also need the following lemma of Grzesik et al. (2022).
Lemma 7 (cf. Lemma 4.2 of Grzesik et al. (2022)).
@@
Acknowledgements.
We are grateful to Jakub Gajarský, Lars Jaffke, Paloma Thomé de Lima, Jana Masařiková, Uéverton Souza, and Michał Pilipczuk for inspiring discussions and to Martin Milanič for bringing our attention to Open problem 3 of  Milanič and Pivač (2021).